# Constrained Decoding for Robotics Foundation Models constrained-robot-fms.github.io

### v1

@@ 4.2 Benchmarks and Research Questions
We compare our proposed techniques with (1) an unconstrained base model and (2) a base model with a filtering mechanism. The filtering mechanism picks a default action (turning left or right in place) upon predicted violation of the safety specification, similar to the Simplex architecture [23]. Simplex architecture is a classic scheme in which a high-performance advanced controller is continuously monitored by a provably safe but less capable backup controller. Simplex based techniques have been used extensively for safety-critical robotics and are a widely accepted standard for runtime-safety comparisons. We are interested in three main metrics: STL Satisfaction Rate (STL St): Proportion of trajectories that satisfy the specified STL formula, Task Success rate (SR): Standard task success, Average Robustness Score (AR): A quantitative measure of how strongly STL constraints are satisfied. The three main research questions we investigated in this paper:
@@
    RQ1: Do RCD and HCD outperform the baselines in STL satisfaction?
@@
    RQ2: Do RCD and HCD lead to task success rates comparable to those of the unconstrained technique?
@@
    RQ3: Does RCD perform better than HCD for task success rates?
Figure 2: Distribution of STL robustness scores across three action selection techniques for $\phi_{avoid}$. The red dashed line marks the violation threshold at robustness = 0. HCD achieves the lowest violation rate, while RCD strikes a balance between task success and STL satisfaction.
Specification Decoding STL St (%) SR (%) AR
$\phi_{avoid}$ Unconstrained (SPOC) 72.00 82.5 1.60
Filtering (SPOC-F) 97.00 72.00 1.57
Hard-Constrained (SPOC-HCD) 97.00 72.50 1.73
Robust-Constrained (SPOC-RCD) 93.00 76.00 1.71
$\phi_{geofence}$ Unconstrained (SPOC) 78.00 81.5 1.68
Filtering (SPOC-F) 100.00 72.00 1.89
Hard-Constrained (SPOC-HCD) 100.00 76.5 1.97
Robust-Constrained (SPOC-RCD) 95.35 80.23 1.86
Table 1: Comparison of decoding strategies across different backbone models on STL-constrained tasks. Our constrained decoding approaches show significant improvements in STL satisfaction and robustness while preserving task success.
@@ 4.3 Results
Our results are highlighted in Table 1 and Figure 2. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
@@
[…]
@@
While both HCD and RCD improve safety over the unconstrained baseline, they differ in how they balance constraint satisfaction with task success. HCD enforces strict STL satisfaction that results in frequent conservatism and lower successful task completion rates. In contrast, RCD’s soft penalization leads to higher task success—up to 4–5 % more than HCD—while still maintaining over 93 % STL satisfaction. These results show that RCD achieves a better trade-off between safety and goal-directed behavior, especially in settings where occasional low-risk actions can lead to higher long-term rewards. We also perform an ablation for varying values of $\beta$ to investigate this tradeoff and provide these results in Appendix C.
Our proposed techniques effectively enforces STL specifications during policy execution. HCD ensures full compliance but occasionally sacrifices task success due to strict truncation. RCD strikes a balance, offering high satisfaction rates and robust performance. This highlights the feasibility of combining learning-based models with formal safety constraints.
@@ 5 Related Work
Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) [24], safe reinforcement learning [25], and temporal logic-based shielding approaches [26]. Recently, with the advent of Vision Language Action models and their impressive generalizable capabilities for manipulation, navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models. Although classical methods offer formal guarantees, they require pretraining/fine-tuning stage interventions, which can be restrictive. For example, SafeVLA [27] fine-tunes pre-trained foundation models with task-specific safety costs, achieving strong performance in Safety-CHORES tasks. However, the safety specification is expected to be embedded in the training data and loss, meaning the model cannot generalize to new safety constraints at test time. In contrast, ASIMOV [28] explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees. Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model and incurring negligible computational overhead. The closest to our work is SELP [29] that proposes LTL-constrained decoding for language model-based plan generation. However, since LTL does not possess quantiative semantics, they are restricted to hard masking-based approach.
@@ 6 Limitations and Future Work
In this work, we introduce a constrained decoding framework for enforcing STL specifications for robotic foundation models. Our approach enables runtime adaptation to novel safety specifications without retraining. Through experiments across multiple simulated environments, we demonstrated that our method significantly improves STL satisfaction while maintaining high task success rates. Our approach makes two critical assumptions that can be a limiting factor. First, we assume access to specifications that are defined over the state space and that these specifications are generated by roboticists. Although this is a common situation for safety critical deployment contexts like aerial robotics [30, 31], these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation. We hope to remedy this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic (ETL) [32] and using Large Language Models for generating high level specifications automatically [33]. Second, our approach also assumes access to a predictive model to evaluate the impact of actions on future trajectories. This can limit applicability to settings where accurate dynamics models are unavailable. However, it is possible to mitigate this via learned dynamics models [34] or world models proposed in [35, 36]. In addition, while our experiments are conducted in simulation, the underlying foundation model (SPOC) has been successfully deployed on real robots with minimal sim-to-real degradation [2]. Since our constrained decoding operates primarily at inference time and leaves the base model unchanged, the observed improvements in STL satisfaction and task success are expected to transfer over to the real world. We plan to rigorously test this by deploying SpecDec on the same physical robot platform as [2].
Appendix
Appendix A Quantiative Semantics of STL
@@
Appendix B HCD proof
We rewrite the theorem here for easier comprehension.
Let $\pi_{\text{HCD}}$ denote the policy induced by hard-constrained decoding with respect to an STL specification $\varphi$, and let $f$ be a deterministic dynamics model used during decoding. Then the generated trajectory satisfies the specification:
$\{(x_{t+1},\hat{a}_{t+1}),\dots,(x_{T},\hat{a}_{T})\}\models\varphi$
where each action $\hat{a}_{t+k}\sim\pi_{\text{HCD}}(\cdot\mid x_{t+k-1})$, and $x_{t+k}=f(x_{t+k-1},\hat{a}_{t+k})\quad\text{for all }k.$
Proof Sketch.
[…]
We proceed by induction over time. At step $t+1$, the masking ensures that the selected action $a_{t+1}$ is such that $(x_{t+1},a_{t+1})\models\varphi$, where $x_{t+1}=f(x_{t},a_{t+1})$. Assuming that the sequence $\{(x_{t+1},a_{t+1}),\dots,(x_{t+k-1},a_{t+k-1})\}$ satisfies $\varphi$, we show that $(x_{t+k},a_{t+k})$ will also satisfy $\varphi$. Since actions are masked based on predicted rollouts using $f$, and $f$ is exact, the only permitted actions at step $t+k$ are those that extend the sequence without violating $\varphi$. By induction, the entire sequence satisfies $\varphi$. ∎
Appendix C Robustness weighting ablation
Figure 3: Success rate and STL satisfaction rate across different values of $\beta$ for RCD, which controls the importance of STL spcification during decoding. As $\beta$ increases, STL satisfaction improves while task success shows a slight tradeoff beyond moderate values.
To evaluate the impact of relative weighting between robustness and base logits, we plotted the success rate and STL satisfaction rate across varying values for $\beta$. As shown in Figure 3, increasing the robustness threshold initially improves both task success and specification satisfaction, with success rates peaking around a moderate threshold (5–10). However, at higher thresholds (20 and 25), success rates begin to decline slightly, while STL satisfaction remains high.
This trend highlights a trade-off introduced by weighting robustness. Higher robustness weighting leads to better STL satisfaction at the cost of hindering overall task completion as logits for feasible actions are shifted heavily by the robustness values. These results demonstrate the importance of carefully balancing robustness weighting to maintain both task success and STL satisfaction.

### v3

@@ 4.2 Experimental setup
We compare our proposed techniques with (1) an unconstrained base model and (2) a base model with a filtering mechanism. The filtering mechanism picks a default action (turning left or right in place) upon predicted violation of the safety specification, similar to the Simplex architecture [29]. Simplex architecture is a classic scheme in which a high-performance advanced controller is continuously monitored by a provably safe but less capable backup controller. Simplex based techniques have been used extensively for safety-critical robotics and are a widely accepted standard for runtime-safety comparisons. We evaluate performance using two main metrics: STL Satisfaction Rate (STL St), defined as the proportion of trajectories that satisfy the specified STL formula, and Task Success Rate (SR), which measures standard task success. The three main research questions we investigated in this paper:
@@
    RQ1: Do HCD and RCD provide higher STL satisfaction than the unconstrained baselines?
@@
    RQ2: Do HCD and RCD preserve task success rates comparable to the unconstrained baselines?
@@
    RQ3: Does RCD achieve better task success than HCD while maintaining high STL satisfaction?
$\bm{\phi}_{avoid}$: STL / SR (% $\uparrow$) $\bm{\phi}_{geofence}$: STL / SR (% $\uparrow$)
Decoding SPOC Flare PoliFormer SPOC Flare PoliFormer
Unconstrained 72.0 / 82.5 75.5 / 82.0 77.0 / 82.5 78.0 / 81.5 68.0 / 81.0 73.0 / 81.5
Filtering 100.0 / 72.0 100.0 / 78.5 100.0 / 75.5 100.0 / 72.0 100.0 / 66.5 100.0 / 67.5
HCD 100.0 / 72.5 100.0 / 81.0 100.0 / 78.5 100.0 / 76.5 100.0 / 67.5 100.0 / 72.5
RCD 93.0 / 76.0 83.0 / 82.5 87.5 / 83.5 95.5 / 80.0 80.0 / 71.5 85.5 / 77.5
Table 1: Comparison by decoding technique across models (SPOC, FLARE, PoliFormer) for specifications $\phi_{avoid}$ and $\phi_{geofence}$. Each cell reports STL satisfaction / success rate (%). Higher is better (↑).
(a) Unconstrained v/s HCD
(b) Unconstrained v/s RCD
Figure 2: Qualitative comparison of decoded trajectories for a sample scene. Each plot shows a top down view of an overlay of trajectories starting from the white dot under the instruction “find an alarm clock”. The unconstrained model passes through two forbidden regions (red squares) on the way to the target object located on the table. In contrast, HCD (left) and RCD (right) modify the trajectories to respect STL safety specifications while still reaching the goal. More visualizations are available in the Appendix A.2.
@@ 4.3 Results
Our results are highlighted in Table 1. We also visualize sample trajectories in Figure 2 for one scene and task. Unless stated otherwise, all numbers are averaged over $200$ evaluation episodes.
@@
Simplex-style filtering attains high satisfaction but sacrifices task success because the agent takes predefined safe actions. HCD shows similar behavior: although safety is maximized, success rates are consistently 5–10% lower than the baseline across models and specifications. However, as HCD factors in base model logits, it is able to achieve higher task satisfaction compared to Simplex-style filtering. In contrast, RCD preserves success rates much closer to the unconstrained level. For $\phi_{avoid}$, RCD achieves 82–85% success compared to 82–83% for the unconstrained controllers for Flare and PoliFormer. For $\phi_{geofence}$, it maintains 77–80% compared to 81–82% unconstrained for SPOC and PoliFormer. However, we note that RCD does not fully recover success in every case: on SPOC with $\phi_{avoid}$ and Flare with $\phi_{geofence}$, task success remains several points below the unconstrained baseline. Nevertheless, RCD enforces safety while avoiding the large performance penalty observed with filtering.
@@
While both HCD and RCD improve safety over the unconstrained baseline, they differ in how they balance constraint satisfaction with task success. HCD enforces strict STL satisfaction that results in frequent conservatism and lower successful task completion rates. In contrast, RCD’s soft penalization leads to higher task success while still maintaining reasonable STL satisfaction. These results show that RCD achieves a better trade-off between safety and goal-directed behavior, especially in settings where occasional low-risk actions can lead to higher long-term rewards.
Overall, Our proposed techniques effectively enforce safety STL specifications during policy execution. HCD ensures full compliance, but occasionally sacrifices task success due to strict truncation. RCD strikes a balance, offering high satisfaction rates and robust performance. This highlights the feasibility of combining learning-based models with formal safety constraints.
4.4 Ablation studies
Does inaccurate dynamics modeling substantially reduce STL satisfaction? In this work, we assume a simple unicycle dynamics model due it’s generalization capability for diverse robotic platforms. Although this represents a high-level abstraction of true dynamics, such modeling simplifications are standard in the formally assured robot safety literature [28]. However, both RCD and HCD depend on this assumption and inaccurate modeling can impact STL satisfaction. To evaluate the impact of inaccurate dynamics modeling, we conducted an ablation in which we inject gaussian perturbations into the dynamics (0.01 m per step translational noise i.e. 5% of nominal forward step, 1 ° per step rotational noise i.e. 3.3% of yaw step) for both HCD and RCD. Our results are visualized in Figure 3. Across all three base models, the drop in STL satisfaction rates from baseline to noisy dynamics is relatively small. We observe that SafeDec shows graceful degradation under significant per-step disturbances.
(a) Ablation with noisy dynamics
(b) Ablation on varying values of $\beta$ for RCD
Figure 3: Ablation studies. (a) STL satisfaction (%) for HCD and RCD under baseline vs noisy dynamics across base models. (b) Effect of $\beta$ on success rate and safety satisfaction.
How do varying values of $\beta$ affect success rate and STL satisfaction in RCD? As highlighted in section 3.3, RCD uses a hyperparameter $\beta$ that modulates the trade-off between specification adherence and the original task objective. The value of $\beta$ affects the relative weighting between robustness and base logits. To investigate the impact of $\beta$ on the success rate and STL satisfaction, we performed an ablation with varying values for $\beta$ for Flare and PoliFormer. We observe that as $\beta$ increases for PoliFormer, both STL satisfaction and success rate improve in tandem until $\beta$ = 10, suggesting that moderate regularization can actually aid policy execution. Beyond this, STL satisfaction continues to improve but at the cost of lower success rates. For Flare, larger $\beta$ values improve STL satisfaction but reduce success rates. These results highlight that the influence of $\beta$ is model-dependent but in general demonstrate that SafeDec provides a tunable mechanism to balance safety and performance objectives.
@@ 5 Related Work
Constraint satisfaction for robotics has been an active area of research that involves techniques such as control barrier functions (CBFs) [30], safe reinforcement learning [31], temporal logic-based shielding and synthesis approaches [32, 33, 34]. Recently, with the advent of vision language action models and their impressive generalizable capabilities for manipulation, navigation and other tasks, there are growing concerns about ensuring safety and correctness without retraining these large models. Although classical methods offer formal guarantees, they either require pretraining/fine-tuning stage interventions or designing a new classical controller for each safety specification, which can be restrictive. For example, SafeVLA [35] fine-tunes pre-trained foundation models with task-specific safety costs, achieving strong performance in Safety-CHORES tasks. However, the safety specification is expected to be embedded in the training data and loss, meaning the model cannot generalize to new safety constraints at test time. In contrast, ASIMOV [36] explores rewriting dangerous instructions with better human-aligned alternatives to steer model behavior without modifying model parameters, but lacks trajectory-level formal guarantees. Our technique achieves a middle ground with the ability to adapt to novel specifications at test time without modifying model parameters while requiring minimal assumptions about the underlying model. The closest to our work is SELP [37] that proposes LTL-constrained decoding for language model-based plan generation. However, SELP is unsuitable for STL because its Boolean predicate-based LTL cannot encode numeric bounds (e.g. $x–x_{goal}||<0.1m$) and does not possess quantitative semantics, which is crucial for ranking actions. These techniques are also tailored to high-level plans from LLMs, not to the per-step low-level actions generated online by policies.
@@ 6 Limitations and Future Work
In this work, we introduce a constrained decoding framework for enforcing safety specifications for large transformer based robot policies. Our approach enables runtime adaptation to novel safety specifications without retraining. Through experiments across multiple simulated environments, we demonstrated that our method significantly improves STL satisfaction while maintaining high task success rates. Our approach makes two critical assumptions that can be a limiting factor. First, we assume access to specifications that are defined over the state space and that these specifications are generated by roboticists. Although this is a common situation for safety critical deployment contexts like aerial robotics [38, 39], these specifications can be difficult to design and involve access to a localization module that can provide accurate state estimation. We plan to overcome this bottleneck by leveraging open-world safety specifications using recent work on embedding spaces-based logic [40] and using large language models for generating high level specifications automatically [41]. Second, our approach also assumes access to an approximate dynamics model to evaluate the impact of actions on future trajectories. While a common assumption for provably safe robotics [28], This can limit applicability of our approach. However, it is possible to mitigate this via learned dynamics models such as MoSim [42] or world models proposed in Zhou et al. [43], Micheli et al. [44] and we plan to explore this in future work.
Appendix A Appendix
A.1 Quantitative Semantics of STL
@@
A.2 Visualizations
To complement our quantitative results, we provide numerous trajectory plots from evaluations across a diverse set of procedurally generated indoor environments. Figure 4 illustrates representative top-down visualizations of trajectories induced by SafeDec .Each plot shows the agent’s starting point and the resulting path under constrained decoding, with red circles marking target objects, green boxes denoting forbidden regions, and orange paths depicting the safe trajectories generated by our method. Together with our quantitative analysis, these qualitative results illustrate the performance of SafeDec across environments and tasks.
Figure 4: Top-down views across indoor environments with SafeDec induced trajectories Red circles mark target objects, green boxes are avoid regions, and orange paths show safe trajectories under constrained decoding.